# Protecting Federated Learning from Extreme Model Poisoning Attacks via Multidimensional Time Series Anomaly Detection

### v2

@@ 5.3 Computational Complexity Analysis
6 EXPERIMENTS
6.1 Experimental Setup
@@
The full details are available in Appendix F.
FL Simulation Environment. To simulate a realistic FL environment, we integrate FLANDERS into Flower [8]. Moreover, we implement in Flower any defense baseline considered that the framework does not natively provide. We set the number $K$ of FL clients to $100$, and we assume that the server selects all these clients in every FL round ($m=K$). In Appendix J.5, we test when the server randomly chooses $m<K$ clients. We suppose the attack starts at $t=1$; then, we monitor the performance of the global model until $t=T=50$. Note that we select the malicious clients randomly across all the participating ones; this implies that a single client can alternate between legitimate and malicious behavior over successive FL rounds.
We describe our FL simulation environment in Appendix 13.
@@
6.2 Attacks
We assess the robustness of FLANDERS under the following well-known attacks. For each attack, we vary the number $b=\lceil r*m\rceil,r\in[0,1]$ of malicious clients, where $r=\{0,0.2,0.6,0.8\}$.
@@ 6.2 Attacks
We detail the parameters of the attacks in Appendix H.
6.3 Evaluation
We evaluate four key aspects of FLANDERS. Firstly, we test its ability to detect malicious clients against the best-competing filtering strategy, FLDetector. Secondly, we measure the accuracy improvement of the global model when FLANDERS is paired with “vanilla” FedAvg and the most popular robust aggregation baselines: FedMedian, Trimmed Mean, Multi-Krum, Bulyan, DnC. Thirdly, we analyze the cost-benefit trade-off of FLANDERS. Lastly, we test the robustness of our method against adaptive attacks. The complete experimental settings are illustrated in Appendix I.
Malicious Detection Accuracy. Tab. 1 shows the precision ($P$) and recall ($R$) of FLDetector and FLANDERS in filtering out malicious clients across different datasets and attacks, with $r=0.2$ (20% evil participants in the FL systems). Remarkably, FLANDERS successfully detects all and only malicious clients across every attack setting except for OPT, outperforming its main competitor, FLDetector. This is further confirmed by Tab. 2, which shows that our method generally provides much higher protection than FLDetector when combined with standard FedAvg. Similar findings are observed for different proportions of malicious clients ($r$) and are reported in Appendix J.1.
Table 1: Precision ($P$) and Recall ($R$) of FLDetector and FLANDERS in detecting malicious clients across various datasets and attacks ($r=0.2$; $T=50$ rounds). Due to space limits, the results for $r=0.6$ and $0.8$ are provided in Tab. 10 and 11 (Appendix J.1).
@@ 6.3 Evaluation
Table 2: Accuracy of the global model using FedAvg with FLDetector and FLANDERS ($r=0.2$). Due to space limits, the results for $r=0.6$ and $0.8$ are provided in Tab. 12 and 13 (Appendix J.1).
Attack FedAvg FLDetector + FedAvg FLANDERS + FedAvg
@@
Aggregation Robustness Lift. We proceed to evaluate the enhancement that FLANDERS provides to the robustness of the global model. Specifically, we measure the best accuracy of the global model under several attack strengths using all the baselines without and with FLANDERS as a pre-filtering strategy. Tab. 3 shows that FLANDERS keeps high accuracy for the best global model under extreme attack ($r=0.8$) when used before every aggregation method, including Multi-Krum and Bulyan, which would otherwise be inapplicable in such strong attack scenarios. This is further emphasized in Tab. 20 (see Appendix J.2), where Multi-Krum, when paired with FLANDERS, can effectively operate without any performance degradation, even in the presence of $80\%$ malicious clients.
Table 3: Accuracy of the global model using all the baseline aggregations without and with FLANDERS ($r=0.8$). Due to space limits, the results for CIFAR-100 are shown in Tab. 14 (Appendix J.2).
Aggregation MNIST Fashion-MNIST CIFAR-10
@@
FedAvg $0.18$ $0.11$ $0.21$ $0.11$ $0.24$ $0.10$ $0.19$ $0.10$ $0.10$ $0.10$ $0.11$ $0.10$
+ FLANDERS ${\bf 0.75}$ ${\bf 0.84}$ ${\bf 0.84}$ ${\bf 0.82}$ ${\bf 0.68}$ ${\bf 0.70}$ ${\bf 0.66}$ ${\bf 0.66}$ ${\bf 0.33}$ ${\bf 0.32}$ ${\bf 0.32}$ ${\bf 0.32}$
FedMedian $0.34$ $0.19$ $0.13$ $0.23$ $0.29$ $0.10$ $0.17$ $0.10$ $0.13$ $0.10$ $0.10$ $0.12$
+ FLANDERS ${\bf 0.81}$ ${\bf 0.84}$ ${\bf 0.85}$ ${\bf 0.81}$ ${\bf 0.70}$ ${\bf 0.71}$ ${\bf 0.71}$ ${\bf 0.68}$ ${\bf 0.29}$ ${\bf 0.29}$ ${\bf 0.31}$ ${\bf 0.28}$
TrimmedMean $0.15$ $0.13$ $0.20$ $0.18$ $0.21$ $0.10$ $0.23$ $0.10$ $0.10$ $0.10$ $0.10$ $0.10$
+ FLANDERS ${\bf 0.81}$ ${\bf 0.83}$ ${\bf 0.83}$ ${\bf 0.85}$ ${\bf 0.71}$ ${\bf 0.70}$ ${\bf 0.71}$ ${\bf 0.69}$ ${\bf 0.30}$ ${\bf 0.29}$ ${\bf 0.30}$ ${\bf 0.29}$
Multi-Krum $0.80$ $0.12$ $0.17$ $0.25$ $0.64$ $0.10$ $0.20$ $0.10$ $0.34$ $0.10$ $0.10$ $0.10$
+ FLANDERS ${\bf 0.87}$ ${\bf 0.90}$ ${\bf 0.88}$ ${\bf 0.89}$ ${\bf 0.69}$ ${\bf 0.68}$ ${\bf 0.72}$ ${\bf 0.68}$ ${\bf 0.38}$ ${\bf 0.38}$ ${\bf 0.39}$ ${\bf 0.40}$
Bulyan N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
+ FLANDERS ${\bf 0.89}$ ${\bf 0.85}$ ${\bf 0.88}$ ${\bf 0.82}$ ${\bf 0.68}$ ${\bf 0.66}$ $0{\bf.68}$ ${\bf 0.70}$ ${\bf 0.40}$ ${\bf 0.43}$ ${\bf 0.40}$ ${\bf 0.41}$
DnC $0.21$ $0.11$ $0.17$ $0.11$ $0.25$ $0.10$ $0.14$ $0.10$ $0.10$ $0.10$ $0.10$ $0.10$
+ FLANDERS ${\bf 0.85}$ ${\bf 0.87}$ ${\bf 0.89}$ ${\bf 0.87}$ ${\bf 0.71}$ ${\bf 0.69}$ ${\bf 0.68}$ ${\bf 0.68}$ ${\bf 0.41}$ ${\bf 0.40}$ ${\bf 0.39}$ ${\bf 0.40}$
For weaker attacks, e.g., $r=0.2$, FLANDERS still generally improves the accuracy of the global model, except when combined with Bulyan, which alone appears already robust enough to counter these mild attacks (see Tab. 15 and 16 in Appendix J.2). The complete results are in Appendix J.2.
@@
Cost-Benefit Analysis. Obviously, the robustness guaranteed by FLANDERS under extreme attack scenarios comes with costs, especially due to the MAR estimation stage. Fig. 4 of Appendix J.3 depicts two scatter plots for the MNIST and CIFAR-10 datasets, focusing on a specific attack scenario (AGR-MM). Each data point on a scatter plot represents a method under one of two attack strengths considered ($r=0.2$ and $r=0.6$). These data points are specified by two coordinates: the overall training time on the $x$-axis and the maximum accuracy of the global model on the $y$-axis.
@@
Robustness against Adaptive Attacks. We test the robustness of our method when malicious clients might exploit the knowledge that the FL server implements our FLANDERS filter. We consider two scenarios: one more plausible and the other most pessimistic. For the first setup, Table 21 in Appendix J.4 shows that FLANDERS can cope with adaptive attackers who tentatively guess the subset of parameters used by the FL server to estimate the MAR forecasting model (non-omniscient adaptive attack). In contrast, Table 22 highlights that FLANDERS becomes unsurprisingly ineffective in the worst-case scenario, i.e., if malicious clients know exactly which parameters are used by the FL server (omniscient adaptive attack). However, this latter situation is quite unrealistic in practice.
7 CONCLUSION

### v3

@@ V-E Computational Complexity Analysis
VI Experiments
VI-A Experimental Setup
@@
The full details are available in Appendix XII.
FL Simulation Environment. To simulate a realistic FL environment, we integrate FLANDERS into Flower [8]. Moreover, we implement in Flower any defense baseline considered that the framework does not natively provide. In Table II we synthetically report the hyperparameters.
We set the number $K$ of FL clients to $100$, and we assume that the server selects all these clients in every FL round ($m=K$). In Appendix XV-B, we test when the server randomly chooses $m<K$ clients. We suppose the attack starts at $t=1$; then, we monitor the performance of the global model until $t=T=50$. Recall that we select the malicious clients randomly across all the participating ones; this implies that a single client can alternate between legitimate and malicious behavior over successive FL rounds.
TABLE II: Main properties of our FL environment simulated on Flower.
Total N. of Clients $K=100$
N. of Selected Clients (at each round) $m=K=100$
Ratio of Malicious Clients $r=\{0,0.2,0.6,0.8\}$
Total N. of FL Rounds $T=50$
Autoregressive Order of MAR $w=1$
Historical Window Size of Past FL Rounds $l=2$
Non-IID Dataset Distribution across Clients $\alpha_{D}=0.5$
@@
VI-B Attacks
We assess the robustness of FLANDERS under the following well-known attacks. For each attack, we vary the number $b=\lceil r*m\rceil,r\in[0,1]$ of malicious clients, where $r=\{0,0.2,0.6,0.8\}$. Ultimately, we analyze an attacker’s performance embedded with a MAR model.
@@ VI-B Attacks
Below, we describe the critical parameters for each attack considered, which are also summarized in Table III.
LIE. This method has no parameters to set.
AGR-MM. In addition to the threshold $\tau$, AGR-MM uses the perturbation vectors $\nabla^{p}$ in combination with the scaling coefficient $\gamma$ to optimize. We set $\tau=10^{-5}$ and $\nabla^{p}=\{std\}$, which is the vector obtained by computing the parameters’ inverse of the standard deviation. For Krum, we set $\nabla^{p}=\{uv\}$, which is the inverse unit vector perturbation.
TABLE III: Key parameter settings for each attack strategy considered.
Attack Parameters
GAUSS $\sigma=10$
LIE N/A
OPT $\tau=10^{-5}$
AGR-MM $\tau=10^{-5};\nabla^{p}=\{uv,std\};\gamma=5$
VI-C Evaluation
We evaluate four key aspects of FLANDERS. Firstly, we test its ability to detect malicious clients against the best-competing filtering strategy, FLDetector. Secondly, we measure the accuracy improvement of the global model when FLANDERS is paired with “vanilla” FedAvg and the most popular robust aggregation baselines: FedMedian, Trimmed Mean, Multi-Krum, Bulyan, DnC. Thirdly, we analyze the cost-benefit trade-off of FLANDERS. Lastly, we test the robustness of our method against adaptive attacks.
Malicious Detection Accuracy. Table IV shows the precision ($P$) and recall ($R$) of FLDetector and FLANDERS in filtering out malicious clients across different datasets and attacks, with $r=0.2$ (20% evil participants in the FL systems). Remarkably, FLANDERS successfully detects all and only malicious clients across every attack setting except for OPT, outperforming its main competitor, FLDetector. This is further confirmed by Table V, which shows that our method generally provides much higher protection than FLDetector when combined with standard FedAvg.
Better results are observed for extreme attack settings ($r=0.8$) in Tables VI, and VII, where OPT maximizes the distance between the legitimate models and the malicious ones by erroneously thinking that the high number of its controlled clients pierces through the aggregation functions. Results for $r=0.6$ in Appendix XV-C are consistent with these findings.
FLDetector fails to distinguish malicious updates from legitimate ones because (i) the anomaly score is based on approximating a single Hessian matrix that must be close to all legitimate clients, making it inadequate for highly non-iid settings, where malicious updates can be similar to legitimate ones; and (ii) the suspicious scores are computed as the average normalized Euclidean distance of the past iterations. Although this is coherent and works well with a threat model where the malicious clients are always the same across all rounds, FLANDERS does not make such an assumption, making FLDetector unable to recognize malicious updates arriving for the first time from an ex-legitimate client.
TABLE IV: Precision ($P$) and Recall ($R$) of FLDetector and FLANDERS in detecting malicious clients across various datasets and attacks ($r=0.2$; $T=50$ rounds).
@@ VI-C Evaluation
TABLE V: Accuracy of the global model using FedAvg with FLDetector and FLANDERS ($r=0.2$).
FedAvg FLDetector + FedAvg FLANDERS + FedAvg
@@
TABLE VI: Precision ($P$) and Recall ($R$) of FLDetector and FLANDERS in detecting malicious clients across various datasets and attacks ($r=0.8$; $T=50$ rounds).
FLDetector FLANDERS
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
$P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$ $P$ $R$
TABLE VII: Accuracy of the global model using FedAvg with FLDetector and FLANDERS ($r=0.8$).
FedAvg FLDetector + FedAvg FLANDERS + FedAvg
@@
Aggregation Robustness Lift. We proceed to evaluate the enhancement that FLANDERS provides to the robustness of the global model. Specifically, we measure the best accuracy of the global model under several attack strengths using all the baselines without and with FLANDERS as a pre-filtering strategy. Table VIII shows that FLANDERS keeps high accuracy for the best global model under extreme attacks ($r=0.8$) when used before every aggregation method, including Multi-Krum and Bulyan, which would otherwise be inapplicable in such strong attack scenarios. This is further emphasized in Table XX (see Appendix XV-D), where Multi-Krum, when paired with FLANDERS, can effectively operate without any performance degradation, even in the presence of $80\%$ malicious clients.
In Figure 3 we compare the accuracy of the global model when using FLANDERS (left) and when using FedAvg (right) across the whole FL training process. The left figure shows how the evolution of the accuracy over multiple rounds remains stable and similar to the one without any attack (dashed line), while on the right the accuracy drops irremediably.
Figure 3: FedAvg with FLANDERS (left) vs. ”vanilla” FedAvg (right). Accuracy of the global model in each FL round under all attack strategies on the MNIST dataset, with $80\%$ of malicious clients. Attack starts at round $t=3$.
TABLE VIII: Accuracy of the global model using all the baseline aggregations without and with FLANDERS ($r=0.8$).
MNIST Fashion-MNIST CIFAR-10 CIFAR-100
GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM GAUSS LIE OPT AGR-MM
FedAvg $0.18$ $0.11$ $0.21$ $0.11$ $0.24$ $0.10$ $0.19$ $0.10$ $0.10$ $0.10$ $0.11$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.75}$ ${\bf 0.84}$ ${\bf 0.84}$ ${\bf 0.82}$ ${\bf 0.68}$ ${\bf 0.70}$ ${\bf 0.66}$ ${\bf 0.66}$ ${\bf 0.33}$ ${\bf 0.32}$ ${\bf 0.32}$ ${\bf 0.32}$ ${\bf 0.09}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.10}$
FedMedian $0.34$ $0.19$ $0.13$ $0.23$ $0.29$ $0.10$ $0.17$ $0.10$ $0.13$ $0.10$ $0.10$ $0.12$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.81}$ ${\bf 0.84}$ ${\bf 0.85}$ ${\bf 0.81}$ ${\bf 0.70}$ ${\bf 0.71}$ ${\bf 0.71}$ ${\bf 0.68}$ ${\bf 0.29}$ ${\bf 0.29}$ ${\bf 0.31}$ ${\bf 0.28}$ ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.10}$
TrimmedMean $0.15$ $0.13$ $0.20$ $0.18$ $0.21$ $0.10$ $0.23$ $0.10$ $0.10$ $0.10$ $0.10$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.81}$ ${\bf 0.83}$ ${\bf 0.83}$ ${\bf 0.85}$ ${\bf 0.71}$ ${\bf 0.70}$ ${\bf 0.71}$ ${\bf 0.69}$ ${\bf 0.30}$ ${\bf 0.29}$ ${\bf 0.30}$ ${\bf 0.29}$ ${\bf 0.11}$ ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.11}$
Multi-Krum $0.80$ $0.12$ $0.17$ $0.25$ $0.64$ $0.10$ $0.20$ $0.10$ $0.34$ $0.10$ $0.10$ $0.10$ $0.08$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.87}$ ${\bf 0.90}$ ${\bf 0.88}$ ${\bf 0.89}$ ${\bf 0.69}$ ${\bf 0.68}$ ${\bf 0.72}$ ${\bf 0.68}$ ${\bf 0.38}$ ${\bf 0.38}$ ${\bf 0.39}$ ${\bf 0.40}$ ${\bf 0.11}$ ${\bf 0.10}$ ${\bf 0.10}$ ${\bf 0.11}$
Bulyan N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
+ FLANDERS ${\bf 0.89}$ ${\bf 0.85}$ ${\bf 0.88}$ ${\bf 0.82}$ ${\bf 0.68}$ ${\bf 0.66}$ ${\bf 0.68}$ ${\bf 0.70}$ ${\bf 0.40}$ ${\bf 0.43}$ ${\bf 0.40}$ ${\bf 0.41}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.11}$
DnC $0.21$ $0.11$ $0.17$ $0.11$ $0.25$ $0.10$ $0.14$ $0.10$ $0.10$ $0.10$ $0.10$ $0.10$ $0.01$ $0.01$ $0.01$ $0.01$
+ FLANDERS ${\bf 0.85}$ ${\bf 0.87}$ ${\bf 0.89}$ ${\bf 0.87}$ ${\bf 0.71}$ ${\bf 0.69}$ ${\bf 0.68}$ ${\bf 0.68}$ ${\bf 0.41}$ ${\bf 0.40}$ ${\bf 0.39}$ ${\bf 0.40}$ ${\bf 0.10}$ ${\bf 0.11}$ ${\bf 0.11}$ ${\bf 0.12}$
For weaker attacks, e.g., $r=0.2$, FLANDERS still generally improves the accuracy of the global model, except when combined with Bulyan, which alone appears already robust enough to counter these mild attacks (see Table XVIII in Appendix XV-D). The complete results are in Appendix XV-D. However, it is worth remarking that current robust aggregation methods like Multi-Krum and Bulyan are not designed to handle stronger attack settings. Nevertheless, when these methods are combined with FLANDERS, they can be deployed successfully even under extremely severe attacks due to FLANDERS’ capability to filter out every malicious client before the aggregation process takes place. This is illustrated with an example in Table XX, where it is demonstrated that Multi-Krum, when paired with FLANDERS, can effectively operate without any performance degradation even when the proportion of malicious clients reaches $80\%$. Similarly, as shown in Fig. 3, the performance remains robust across all FL rounds even when FLANDERS is paired with a simple aggregation function, such as FedAvg. Instead, when our filter is removed, the accuracy drops as soon as the attack begins.
@@
Cost-Benefit Analysis. Obviously, the robustness guaranteed by FLANDERS under extreme attack scenarios comes with costs, especially due to the MAR estimation stage. Fig. 4 depicts two scatter plots for the MNIST and CIFAR-10 datasets, focusing on a specific attack scenario (AGR-MM). Each data point on a scatter plot represents a method under one of two attack strengths considered ($r=0.2$ and $r=0.6$). These data points are specified by two coordinates: the overall training time on the $x$-axis and the maximum accuracy of the global model obtained after $T=50$ FL rounds on the $y$-axis.
Figure 4: Accuracy vs. total Training Time (in seconds) of FedAvg and Bulyan compared with their corresponding versions with FLANDERS as a filter for the MNIST (left) and CIFAR-10 (right) datasets.
@@
Robustness against Adaptive Attacks. In this section, we further validate the robustness of FLANDERS against adaptive attacks. We consider a scenario where malicious clients are aware that the FL server uses our method as a pre-aggregation filter. Specifically, we focus on two levels of knowledge. The first scenario assumes that malicious clients tentatively guess the subset of parameters ($\widetilde{d}$) used by the FL server to estimate the MAR forecasting model. We refer to this setting as non-omniscient. The second, more challenging as well as unrealistic scenario assumes that malicious clients know exactly which parameters are used by the FL server. We call this second scenario omniscient. Obviously, the latter penalizes FLANDERS way more than the former.
As introduced above, we first assume that the attacker is non-omniscient, meaning it does not know which parameters the server has selected for the MAR estimation. Instead, the attacker selects the last layer as $\widetilde{d}$. Afterward, we consider an omniscient attacker who exploits the knowledge of the parameters selected by the server. In Table IX, we show the results of the non-omniscient scenario, where FLANDERS + Multi-Krum outperforms all other baselines on all three datasets. Table X, on the other hand, refers to the omniscient scenario and demonstrates a different pattern, where FedAvg and Multi-Krum alone perform better than when coupled with FLANDERS. This may be because FLANDERS consistently filters out legitimate local models in favor of corrupted ones. When using FedAvg, the impact of corrupted parameters is mitigated by averaging a larger number of legitimate models and because the corrupted models’ parameter values are not too different, unlike in methods like OPT. On the other hand, Multi-Krum selects parameters with more nearby neighbors, and with only $b=5$ corrupted clients, legitimate models likely still have more and closer neighbors, effectively defending against our adaptive attack.
TABLE IX: Accuracy of the global model using FedAvg and Multi-Krum, with and without FLANDERS, under the non-omniscient adaptive attack.
Strategy MNIST Fashion-MNIST CIFAR-10
FedAvg $0.84$ $0.68$ $0.45$
FLANDERS + FedAvg $0.82$ $0.67$ $0.43$
Multi-Krum $0.87$ ${\bf 0.72}$ $0.41$
FLANDERS + Multi-Krum ${\bf 0.90}$ ${\bf 0.72}$ ${\bf 0.47}$
TABLE X: Accuracy of the global model using FedAvg and Multi-Krum, with and without FLANDERS, under the omniscient adaptive attack.
Strategy MNIST Fashion-MNIST CIFAR-10
FedAvg $0.78$ ${\bf 0.68}$ ${\bf 0.25}$
FLANDERS + FedAvg $0.65$ $0.61$ $0.10$
Multi-Krum ${\bf 0.86}$ ${\bf 0.68}$ $0.23$
FLANDERS + Multi-Krum $0.73$ $0.60$ $0.10$
VII Limitations and Future Work
VII-A Efficiency/Feasibility
As we discussed in Section 5.3 and Appendix XIV, FLANDERS may suffer from a high computational cost that could limit its deployment in practice. This concern holds particularly true for cross-device FL configurations encompassing millions of edge devices. Conversely, the impact on cross-silo FL scenarios would be notably less pronounced. However, as we have introduced FLANDERS as a versatile and robust aggregation approach applicable to diverse FL setups (cross-silo and cross-device), there are implementation techniques available to mitigate its complexity. For instance, methods like random parameter sampling [37] can be employed, and we have already incorporated them. Still, we plan to enhance the scalability of FLANDERS further in future work. For example, we could replace the standard matrix inversion algorithm with the more efficient Coppersmith-Winograd algorithm [16] and find the optimal frequency for re-estimating FLANDERS’ parameters, instead of performing it during each FL round.
In an extended version of this work, we plan to include a parameter sensitivity analysis, where the performance of FLANDERS is evaluated based on the number $d$ of sampled parameters. This analysis will hopefully provide insights into the optimal trade-off between robustness and efficiency. Additionally, more sophisticated parameter selection strategies beyond pure random sampling could be explored to focus on the most informative neurons, such as Neuron Shapley [22].
VII-B Potential Privacy Leakage
VII-C Cross-Device Setting
The cross-device setting (thousands to millions of clients) penalizes FLANDERS, as the server cannot select them all, and the probability of choosing the same client in consecutive rounds is low. However, FLANDERS is more appropriate in cross-silo FL (tens to hundreds of clients) than in cross-device settings. In fact, large attacks involving more than 50% of clients are less feasible when the total number of participants grows to the order of thousands [38]. In such cases, since selected clients have little or no history, FLANDERS computes distances between local updates and the last global model, turning it into a heuristic similar to (Multi-)Krum. Thus, FLANDERS will be comparable to any other heuristic.
VII-D Benchmarking
In addition, robust federated aggregation is a hot research topic; keeping pace with the massive body of work that has been flourishing is challenging. Hence, we might have missed considering some other methods. However, we believe that the value of our work still stands. In this regard, our contribution is clear: We are the first to frame the problem of detecting untargeted model poisoning attacks on FL as a matrix-valued time series anomaly detection task and to propose a method effective under severe attack settings, as opposed to existing baselines.
VIII Conclusion